OpenStudy (anonymous):

Solve the system of equations for x and y. y = x + 5 2x + 2y = 30 Select one: a. (9, 4) b. (10, 5) c. (4, 9) d. (5, 10)

OpenStudy (anonymous):

2x+2(x+5)=30 2x+2x+10=30 x=5 y=5+5 y=10

OpenStudy (anonymous):

so it D because y is 10

OpenStudy (anonymous):

Yep!